My head is spinning I really don't know this please show how thank you so much.

Set K contains 50 elements, set J contains 66 elements, and 14 elements are common to both sets. Find n (K U J).

Answer by stanbon(75887)

Set K contains 50 elements, set J contains 66 elements, and 14 elements are common to both sets. Find n (K U J).
--------------------
Draw two intersecting circles.
Label them K and J.
---------------
Put "14" in the intersection because 14 elements are common to both sets.
----------
Then the rest of K must have 50-14 = 36 elements.
------------
And the rest of J must have 66-14 = 52 elements.
---------
So the total number of elements is 52 + 14 + 36 = 102
======================================================
Cheers,
Stan H.